# Mechanistically Interpreting a Transformer-based 2-SAT Solver: An Axiomatic Approach

### v1

@@ B.2 Second Layer as Evaluator
We showed in Section 5.1 that the behavior of the first layer is accurately captured by an interpretation which, after the application of $\gamma_{1}$, outputs fixed representative embeddings for each clause at the second literal positions and is constant at all other positions. Prefix replaceability, in particular, enforces this property directly. Furthermore, we must only study attention with the readout token as the destination to explain the output behavior; hence, we can characterize the behavior of attention in the second layer solely by the preferences of each head in the readout token position to the representative clause embeddings in the second literal positions, dramatically reducing dimensionality.
In this way, the key behavior of attention in the second layer is fully described by Figure 9. Each of the four charts shows the preferences on each clause by the corresponding head; the columns correspond to the first literal of the clause and the rows to the second.
@@
While we might read far more into these patterns, this is a mistake in this case. In this instance, the behavior of attention is, in fact, not relevant to the underlying algorithm and serves only to obscure the underlying behavior. As discussed in Section 5.2, we can much better understand attention in concert with the first layer of the MLP; collectively, these components encode the observed evaluating behavior, as discussed in more detail below.
@@
Figure 10: Effect of the residual connection from the second layer’s attention mechanism on the SAT logit. x-axis shows the effect on the SAT logit and y-axis is the number of instances in the analysis test dataset.
@@
Next, we show that the key pathway that determines classification passes through the second layer’s MLP. For an input formula, the effect of the second layer’s attention mechanism through the residual connection on the SAT logit simply the dot product of the post-attention embedding in the readout token position with $W_{U}^{\text{SAT}}$; as we can see in Figure 11, this value is fairly consistent across the dataset and always negative; so a positive SAT logit and hence a SAT classification must depend on the action of the MLP.
@@
As discussed in Section 5.2, only 34 hidden neurons have a significant effect on classification; hence, we focus on the behavior of these neurons. As observed in Figure 11, the effect of the residual stream from the second layer attention on the SAT logit reduces the SAT logit by less than 6 units for all formulas in the analysis test dataset. Furthermore, as observed in Section 5.2, the model always predicts SAT or UNSAT, and the SAT and UNSAT logits are tied (in particular, they are negatives). Hence, the model will predict SAT when the collective effect of these 34 neurons is to increase SAT by > 6 units. We have observed that the expected behavior of these neurons is strongly dependent on which assignments to the variables satisfy the formula $\phi$.
A natural question to ask is whether, whether these neurons will collectively increase the SAT logit sufficiently to output SAT on SAT formulas regardless of which assignment to the variables satisfies the formulas. Calculating this in the same way as we do for an individual neuron in Figure 5, we obtain Figure 11, which shows that, indeed, the collective effect of the evaluating neurons (recall that we refer to the 34 relevant neurons as evaluating neurons) correctly identifies SAT formulas regardless of how the formulas are satisfiable. It may seem odd that the average effect of SAT is significantly below the minimum of the effects for formulas satisfiable with any given assignment, as any SAT formula belongs to one such category; this is an artefact of the fact that no neuron has a large effect for all SAT formulas, and that each specializes in a subset.
@@
However, in reality, individual neurons do not consistently reach activations of $\geq 2$ (for example, see Figure 5, in which the average output value for formulas satisfiable with assignment TFFFF is approximately 1.5), and multiple neurons recognize a given SAT formula. Hence, moderate activation may be necessary across several neurons to predict SAT in some cases; this may explain the redundancy in evaluation behavior between neurons seen in Appendix C. For simplicity in our interpretation (in particular, to allow replacing a numerical weighted sum with Boolean operations) we handle this by using varying thresholds between the corresponding $\alpha$ and $\gamma$ functions for the MLP’s hidden layer (in particular, $\gamma_{2}$ outputs an activation of 2 for neurons whose interpretation predicts high activation, while 0.5 is considered a high activation in the case of $\alpha_{2}$); as noted in Section 5.2.1, if $\gamma_{2}$ and $\alpha_{2}$ both use the lower threshold, replacing the neural network components with our interpretation no longer has a minimal effect on classification.
@@
We’ll focus on the general decision tree classifiers discussed in Section 5.2 here. We train decision tree classifiers on the thresholded activations of each evaluating neuron on the analysis training set; this allows us to learn an arbitrary Boolean function in the features ϕ[]italic-$\phi[...]$. We observe that these decision trees do indeed reliably predict activation. In particular, even very limited-size decision trees are reasonably strong predictors of the behavior of the neurons; we limit our decision trees to four leaf nodes in our experiments. Figures 1314, and 15 illustrate the effect on the decision trees as we lift this constraint. Figure 12 shows that more complex decision trees are not significantly better predictors of the behavior of the neurons.
Figure 13: Decision tree interpretation for unit 29 of the second layer MLP’s hidden layer, maximum leaf nodes: two.
Figure 14: Decision tree interpretation for unit 29 of the second layer MLP’s hidden layer, maximum leaf nodes: four.
Figure 15: Decision tree interpretation for unit 29 of the second layer MLP’s hidden layer, unlimited leaf node count.
@@
To see how the model implements the decision trees, we’ll consider the effect of the composition of the second layer MLP’s hidden layer with the attention mechanism on the representative embeddings of the clauses, using the abstract attention approach discussed earlier. As the pre-softmax attention score by each head on the second literal of a clause is a fixed value for each clause (recall that $\gamma_{1}$ leaves the embedding of the readout token constant), we can express the effect of attention precisely in terms of the number of occurrences of each clause, as we show next.
@@
where $count_{l\lor r}(\phi)$ is the number of occurrences of $l\lor r$ in the formula $\phi$ and where $m$ is the number of attention heads and where $e_{:}$ is the fixed representative embedding of the readout token ‘:’.
@@
Note that the numerator and denominator of the fraction describing the contribution of head $h$ are both linear functions of the number of ocurrences of each clause in the formula.
@@
Listing 2: Second layer’s mechanistic interpretation. First, it applies a series of functions, namely, the neuron interpretations, to the parsed clauses from the first layer (evaluate_satisfiability) and then, it applies an OR operation (predict_satisfiability).
@@
Listing 3: Mechanistic interpretation of full model; see Listings LABEL:fig:layer1_code and LABEL:fig:layer2_code for the component interpretations.
@@
Listing LABEL:fig:layer2_code shows our derived mechanistic interpretation for the second layer. Listing LABEL:fig:abstract_model shows the mechanistic interpretation for the entire model.
Appendix C Neuron Interpretations
@@
Tables 1 and 2 show the decision tree based and disjunction-only neuron interpretations, respectively.
@@ 96 $\phi[TFTTF]$
Appendix D Limitations and Impacts
Limitations.
A limitation of our work is that the proposed axioms are validated by intuitively motivating them and by comparison against past work [40]. We also demonstrate their usage via our case study of mechanistically interpreting a Transformer-based 2-SAT solver. While we believe that our axioms represent a minimal set that it necessary for characterizing a mechanistic interpretation, it may not be sufficient in all cases. In future work, we plan to address this and further validate our axioms by considering analysis of models trained for other problems. We are also able to state Axioms 5 and 6 only in a somewhat informal fashion.
Additionally, while our axioms are well-suited for mechanistic interpretability analyses that aim to understand the model behavior completely, they are not suitable in their present form for circuit analyses that only seek to understand specific aspects of model behavior. We believe that similar axioms can be formulated for circuit analyses but we leave this for future work.
Our techniques for analyzing and interpreting the 2-SAT model require human effort to implement and cannot be automatically applied for the analysis of other models. An exciting yet challenging direction for future research is to develop techniques for automated mechanistic interpretability analysis. We believe that the interpretability techniques that we use and describe (in Section 5 and Appendix B), such as variants of attention pattern analysis and automated learning of decision trees as a component of the mechanistic interpretation, can be generalized to aid in the development of automated interpretability techniques.
Impacts.

### v2

@@ D.2 Second Block as Evaluator
We showed in Section 4.2.1 that the behavior of the first block is accurately captured by an interpretation which, after the application of $\gamma_{1}$, outputs canonical representations for each clause at the second literal positions and is constant at all other positions. Prefix replaceability, in particular, enforces this property directly. Furthermore, we must only study attention with the readout token as the destination to explain the output behavior; hence, we can characterize the behavior of attention in the second block solely by the preferences of each head in the readout token position to the canonical clause representations in the second literal positions, dramatically reducing dimensionality.
In this way, the key behavior of attention in the second block is fully described by Figure 11. Each of the four charts shows the preferences on each clause by the corresponding head; the columns correspond to the first literal of the clause and the rows to the second.
@@
While we might read far more into these patterns, this is a mistake in this case. In this instance, the behavior of attention is, in fact, not relevant to the underlying algorithm and serves only to obscure the underlying behavior. As discussed in Section 4.2.2, we can much better understand attention in concert with the first layer of the MLP; collectively, these components encode the observed evaluating behavior, as discussed in more detail below.
@@
Refer to caption Refer to caption
Figure 12: Effect of the residual connection from the second block’s attention mechanism on the SAT logit. x-axis shows the effect on the SAT logit and y-axis is the number of instances in the analysis test dataset.
@@
Next, we show that the key pathway that determines classification passes through the second block’s MLP. For an input formula, the effect of the second block’s attention mechanism through the residual connection on the SAT logit simply the dot product of the post-attention embedding in the readout token position with $W_{U}^{\text{SAT}}$; as we can see in Figure 13, this value is fairly consistent across the dataset and always negative; so a positive SAT logit and hence a SAT classification must depend on the action of the MLP.
@@
As discussed in Section 4.2.2, only 34 hidden neurons have a significant effect on classification; hence, we focus on the behavior of these neurons. As observed in Figure 13, the effect of the residual stream from the second block attention on the SAT logit reduces the SAT logit by less than 6 units for all formulas in the analysis test dataset. Furthermore, as observed in Section 4.2.2, the model always predicts SAT or UNSAT, and the SAT and UNSAT logits are tied (in particular, they are negatives). Hence, the model will predict SAT when the collective effect of these 34 neurons is to increase SAT by > 6 units. We have observed that the expected behavior of these neurons is strongly dependent on which assignments to the variables satisfy the formula $\phi$.
A natural question to ask is whether, whether these neurons will collectively increase the SAT logit sufficiently to output SAT on SAT formulas regardless of which assignment to the variables satisfies the formulas. Calculating this in the same way as we do for an individual neuron in Figure 3, we obtain Figure 13, which shows that, indeed, the collective effect of the evaluating neurons (recall that we refer to the 34 relevant neurons as evaluating neurons) correctly identifies SAT formulas regardless of how the formulas are satisfiable. It may seem odd that the average effect of SAT is significantly below the minimum of the effects for formulas satisfiable with any given assignment, as any SAT formula belongs to one such category; this is an artifact of the fact that no neuron has a large effect for all SAT formulas, and that each specializes in a subset.
@@
However, in reality, individual neurons do not consistently reach activations of $\geq 2$ (for example, see Figure 3, in which the average output value for formulas satisfiable with assignment TFFFF is approximately 1.5), and multiple neurons recognize a given SAT formula. Hence, moderate activation may be necessary across several neurons to predict SAT in some cases; this may explain the redundancy in evaluation behavior between neurons seen in Appendix G. For simplicity in our interpretation (in particular, to allow replacing a numerical weighted sum with Boolean operations) we handle this by using varying thresholds between the corresponding $\alpha$ and $\gamma$ functions for the MLP’s hidden layer (in particular, $\gamma_{2}$ outputs an activation of 2 for neurons whose interpretation predicts high activation, while 0.5 is considered a high activation in the case of $\alpha_{2}$); as noted in the paragraph titled Hidden Layer of Second Block in Section 4.2.3, if $\gamma_{2}$ and $\alpha_{2}$ both use the lower threshold, replacing the neural network components with our interpretation no longer has a minimal effect on classification.
@@
We’ll focus on the general decision tree classifiers discussed in Section 4.2.2 here. We train decision tree classifiers on the thresholded activations of each evaluating neuron on the analysis training set; this allows us to learn an arbitrary Boolean function in the features ϕ[]italic-$\phi[...]$. We observe that these decision trees do indeed reliably predict activation. In particular, even very limited-size decision trees are reasonably strong predictors of the behavior of the neurons; we limit our decision trees to four leaf nodes in our experiments. Figures 1516, and 17 illustrate the effect on the decision trees as we lift this constraint. Figure 14 shows that more complex decision trees are not significantly better predictors of the behavior of the neurons.
Figure 15: Decision tree interpretation for unit 29 of the second block MLP’s hidden layer, maximum leaf nodes: two.
Figure 16: Decision tree interpretation for unit 29 of the second block MLP’s hidden layer, maximum leaf nodes: four.
Figure 17: Decision tree interpretation for unit 29 of the second block MLP’s hidden layer, unlimited leaf node count.
@@
To see how the model implements the decision trees, we’ll consider the effect of the composition of the second block MLP’s hidden layer with the attention mechanism on the canonical representations of the clauses, using the abstract attention approach discussed earlier. As the pre-softmax attention score by each head on the second literal of a clause is a fixed value for each clause (recall that $\gamma_{1}$ leaves the representation of the readout token constant), we can express the effect of attention precisely in terms of the number of occurrences of each clause, as we show next.
@@
where $count_{l\lor r}(\phi)$ is the number of occurrences of $l\lor r$ in the formula $\phi$ and where $m$ is the number of attention heads and where $e_{:}$ is the fixed representation of the readout token ‘:’.
@@
Note that the numerator and denominator of the fraction describing the contribution of head $h$ are both linear functions of the number of occurrences of each clause in the formula.
@@
Listing 6: Second block’s mechanistic interpretation. First, it applies a series of functions, namely, the neuron interpretations, to the parsed clauses from the first block (evaluate_satisfiability) and then, it applies an OR operation (predict_satisfiability).
@@
Listing 7: Mechanistic interpretation of full model; see Listings LABEL:fig:layer1_code and LABEL:fig:layer2_code for the component interpretations.
@@
Listing LABEL:fig:layer2_code shows our derived mechanistic interpretation for the second block. Listing LABEL:fig:abstract_model shows the mechanistic interpretation for the entire model.
Appendix E Importance of Prefix Axioms: Theoretical Evidence
While it may appear that each componentwise axiom implies the corresponding prefix axiom, for instance, that Axiom 2 implies Axiom 1, this is not the case. While it is possible to derive a bound, the resulting bound is extremely weak, and, for Axioms 3 and 4, we cannot derive even such a weak bound without additional assumptions.
E.1 Equivalence Axioms
Suppose that Axiom 2, component equivalence, holds with a fixed $\epsilon_{0}$. We will show that prefix equivalence (Axiom 1) holds at component $i$ with an $\epsilon$ of $i\epsilon_{0}$; if the number of components $l$ is at least $\lceil{1\over\epsilon_{0}}\rceil$ the bound becomes useless.
Let $PE_{i}$ be the event that $x\sim D$ satisfies the prefix equivalence condition for component $i$, i.e. that $\alpha_{i}\circ d_{t}[:i+1](x)=d_{h}[:i+1]\circ\alpha_{0}(x)$ and and let $CE_{i}$ be the event that $x$ satisfies the component equivalence condition for component $i$, i.e. that $\alpha_{i}\circ d_{t}[:i+1](x)=d_{h}[i]\circ\alpha_{i-1}\circ d_{t}[:i](x)$. Call $Pr[PE_{i}]$ as $p_{i}$.
Now,
$p_{i+1}$ $=\underset{x\sim\mathcal{D}}{Pr}[\alpha_{i+1}\circ d_{t}[:i+2](x)=d_{h}[:i+2]\circ\alpha_{0}(x)]$
$\geq\underset{x\sim\mathcal{D}}{Pr}[\alpha_{i+1}\circ d_{t}[:i+2](x)=d_{h}[i+1]\circ\alpha_{i}\circ d_{t}[:i+1](x)\land d_{h}[i+1]\circ\alpha_{i}\circ d_{t}[:i+1](x)=d_{h}[:i+2]\circ\alpha_{0}(x)]$
$\geq\underset{x\sim\mathcal{D}}{Pr}[\alpha_{i+1}\circ d_{t}[:i+2](x)=d_{h}[i+1]\circ\alpha_{i}\circ d_{t}[:i+1](x)\land\alpha_{i}\circ d_{t}[:i+1](x)=d_{h}[:i+1]\circ\alpha_{0}(x)]$
$=Pr[CE_{i+1}\land PE_{i}],$
noting that $\alpha_{i}\circ d_{t}[:i+1](x)=d_{h}[:i+1]\circ\alpha_{0}(x)$ implies $d_{h}[i+1]\circ\alpha_{i}\circ d_{t}[:i+1](x)=d_{h}[:i+2]\circ\alpha_{0}(x)$ by the definition of the $d_{h}[:]$ notation.
Now, we have that
$Pr[CE_{i+1}\land PE_{i}]$ $=1-Pr[\neg(CE_{i+1}\land PE_{i})]$
$=1-Pr[\neg CE_{i+1}\lor\neg PE_{i}]$
$\geq 1-(Pr[\neg CE_{i+1}]+Pr[\neg PE_{i}]).$
Hence, as $Pr[CE_{i+1}]\geq 1-\epsilon_{0}$ by component equivalence, we have $p_{i+1}\geq 1-(\epsilon_{0}+(1-p_{i}))$ and so $p_{i+1}\geq p_{i}-\epsilon_{0}$. We have $p_{1}\geq 1-\epsilon_{0}$ as prefix equivalence and component equivalence are identical for the first component, and so $p_{i}\geq 1-i\epsilon_{0}$, and we are done.
If we assume that $CE_{i+1}$ and $PE_{i}$ are independent, we could derive a stronger bound:
$Pr[CE_{i+1}\land PE_{i}]=Pr[CE_{i+1}]Pr[PE_{i}]\geq(1-\epsilon_{0})p_{i}$
from which we can derive $p_{i}\geq(1-\epsilon_{0})^{i}$. However, even this stronger bound grows very weak as depth increases, illustrating the need to evaluate a mechanistic interpretation compositionally.
E.2 Replaceability Axioms
We can only derive a bound in the case that $\gamma_{i}=\alpha_{i}^{-1}$ and that all concrete components are invertible. Assume component replaceability, Axiom 4 is satisfied.
Similarly to Appendix E.1 let the $PR_{i}$ and $CR_{i}$ be the events that $x\sim D$ satisfies the prefix replaceability and component replaceability conditions for component $i$, i.e. that $t(x)=d_{t}[i+1:]\circ\gamma_{i}\circ d_{h}[:i+1]\circ\alpha_{0}(x)$ and that $t(x)=d_{t}[i+1:]\circ\gamma_{i}\circ d_{h}[i]\circ\alpha_{i-1}\circ d_{t}[:i](x)$, respectively. By the assumption that all concrete components are invertible, $PR_{i}$ iff $d_{t}[:i+1]=\gamma_{i}\circ d_{h}[:i+1]\circ\alpha_{0}(x)$ and $CR_{i}$ iff $d_{t}[:i+1]=\gamma_{i}\circ d_{h}[i]\circ\alpha_{i-1}\circ d_{t}[:i](x)$. Call $Pr[PR_{i}]$ $p_{i}$.
Now,
$p_{i+1}$ $=\underset{x\sim\mathcal{D}}{Pr}[t(x)=d_{t}[i+2:]\circ\gamma_{i}\circ d_{h}[:i+2]\circ\alpha_{0}(x)]$
$=\underset{x\sim\mathcal{D}}{Pr}[d_{t}[:i+2]=\gamma_{i+1}\circ d_{h}[:i+2]\circ\alpha_{0}(x)]$
$\geq\underset{x\sim\mathcal{D}}{Pr}\left[\begin{array}[]{l}d_{t}[:i+2]=\gamma_{i+1}\circ d_{h}[i+1]\circ\alpha_{i}\circ d_{t}[:i+1](x)\land\\ \gamma_{i+1}\circ d_{h}[:i+2]\circ\alpha_{0}(x)=\gamma_{i+1}\circ d_{h}[i+1]\circ\alpha_{i}\circ d_{t}[:i+1](x)\end{array}\right]$
$\geq\underset{x\sim\mathcal{D}}{Pr}\left[\begin{array}[]{l}d_{t}[:i+2]=\gamma_{i+1}\circ d_{h}[i+1]\circ\alpha_{i}\circ d_{t}[:i+1](x)\land\\ d_{h}[:i+1]\circ\alpha_{0}(x)=\alpha_{i}\circ d_{t}[:i+1](x)\end{array}\right]$
$=\underset{x\sim\mathcal{D}}{Pr}\left[\begin{array}[]{l}d_{t}[i+2:]\circ d_{t}[:i+2]=d_{t}[i+2:]\circ\gamma_{i+1}\circ d_{h}[i+1]\circ\alpha_{i}\circ d_{t}[:i+1](x)\land\\ d_{t}[i+1:]\circ\gamma_{i}\circ d_{h}[:i+1]\circ\alpha_{0}(x)=d_{t}[i+1:]\circ\gamma_{i}\circ\alpha_{i}\circ d_{t}[:i+1](x)\end{array}\right]$
$=\underset{x\sim\mathcal{D}}{Pr}\left[\begin{array}[]{l}t(x)=d_{t}[i+2:]\circ\gamma_{i+1}\circ d_{h}[i+1]\circ\alpha_{i}\circ d_{t}[:i+1](x)\land\\ t(x)=d_{t}[i+1:]\circ\gamma_{i}\circ d_{h}[:i+1]\circ\alpha_{0}(x)\end{array}\right]$
$=Pr[CR_{i+1}\land PR_{i}]$
by the assumption that all concrete components are invertible and that $\gamma_{i}=\alpha_{i}^{-1}$. The proof proceeds as before.
In particular, even with strong conditions such as these, we derive a bound which grows very weak as depth increases, and hence, we observe maintaining Axioms 1 and 3 in addition to Axioms 2 and 4 is essential for a reliable evaluation. See Appendix F for empirical evidence for this fact.
Appendix F Importance of Prefix Axioms: Empirical Evidence
We can demonstrate the importance of the prefix axioms by simulating errors in the interpretation of the 2-SAT model; specifically, every clause output by the first component of the abstract model is replaced by a randomly sampled clause 1% of the time (the slight difference in e.g. the epsilons for component and prefix equivalence for the first component are due to independent sampling); results are shown in Table 1.
The evaluation of the componentwise axioms are identical to the original results except on the first component, as these axioms do not take error propagation into account. The prefix axioms, on the other hand, show the extent to which errors in the earlier components result in downstream errors.
Taking error propagation into account is particularly important when we consider deeper models. While we can derive worst-case compositional bounds from the componentwise results, this may not actually describe model behavior; in particular, as the number of components increases, the derived bound rapidly becomes meaningless (see Appendix E for more details). For example, in our case, the aggregation operation in the final component enables prefix equivalence to improve with depth; hence, restricting ourselves to the pure worst-case analysis derived from component-only analysis would suggest that our interpretation is far less reliable end-to-end than is actually the case.
Table 1: Comparison of results of experiments with synthetic error injected into the first component of the abstract model to original experimental results.
Axiom Noised first abstract component Original abstract model Component 1 Component 2 Component 3 Component 1 Component 2 Component 3 Axiom 1: Prefix Equivalence 0.0955 0.212 0.0315 0.0000374 0.182 0.0128 Axiom 2: Component Equivalence 0.0942 0.182 0.00433 0.0000374 0.182 0.00433 Axiom 3: Prefix Replaceability 0.0583 0.0312 0.0318 0.0418 0.0128 0.0128 Axiom 4: Component Replaceability 0.0581 0.0128 0.00433 0.0418 0.0128 0.00433
Appendix G Neuron Interpretations for the 2-SAT Model
@@
Tables 2 and 3 show the decision tree based and disjunction-only neuron interpretations, respectively.
@@ 96 $\phi[TFTTF]$
Appendix H More Details on Mechanistic Interpretability Analysis of the Modular Addition Model
Pseudocode for the abstract model is shown in Listing LABEL:lst:grokking_code. The concrete model is likewise broken into three corresponding components. The embedding matrix corresponds to encoding_of_inputs, the attention mechanism and the input layer of the MLP correspond to sum_of_angles, and the output layer of the MLP and the unembedding matrix correspond to difference_of_angles_argmax. The abstraction and concretization operators are linear maps learned using the continuous-valued abstract values (e.g. prior to rounding) and the corresponding concrete intermediate representations on the training set, using the original train/test split of Nanda et al. (2023a).
Note that, in Nanda et al. (2023a)’s paper, the third component is further decomposed into two pieces; specifically, the difference-of-angles identities and the summation (along with the argmax) steps are considered separate. The difference-of-angles identity is computed using the MLP output layer and the unembedding matrix while the summation step is computed via the unembedding matrix. As there is no canonical factorization of the unembedding matrix, there exists no natural decomposition of the concrete model into these two components—we analyze the composition of these two components instead.
1modulus = 113
2key_freqs = [14, 35, 41, 42, 52]
3
4cos_sin = tuple[list[float], list[float]]
5
6def encoding_of_inputs(
7 a: int,
8 b: int,
9) -> tuple[cos_sin, cos_sin]:
10 cos_a = [round(cos(2 * f * pi * a / modulus)) for f in key_freqs]
11 cos_b = [round(cos(2 * f * pi * b / modulus)) for f in key_freqs]
12 sin_a = [round(sin(2 * f * pi * a / modulus)) for f in key_freqs]
13 sin_b = [round(sin(2 * f * pi * b / modulus)) for f in key_freqs]
14
15 return (cos_a, sin_a), (cos_b, sin_b)
16
17def sum_of_angles(
18 components: tuple[cos_sin, cos_sin],
19) -> EquivalenceClass:
20 (cos_a, sin_a), (cos_b, sin_b) = components
21 cos_ab = [
22 ca * cb - sa * sb
23 for ca, sa, cb, sb in zip(cos_a, sin_a, cos_b, sin_b)
24 ]
25 sin_ab = [
26 sa * cb + ca * sb
27 for ca, sa, cb, sb in zip(cos_a, sin_a, cos_b, sin_b)
28 ]
29
30 return EquivalenceClass(cos_ab, sin_ab)
31
32def difference_of_angles_argmax(angle_sums_class: EquivalenceClass) -> int:
33 cos_ab, sin_ab = extract_angle_sums(angle_sums_class)
34 cos_ab_minus_c = [
35 [
36 cab * cos(2 * f * pi * c / modulus) + sab * sin(2 * f * pi * c / modulus)
37 for cab, sab, f in zip(cos_ab, sin_ab, key_freqs)
38 ]
39 for c in range(modulus)
40 ]
41
42 return argmax(map(sum, cos_ab_minus_c))
43
44def modular_addition(a: int, b: int) -> int:
45 return difference_of_angles_argmax(
46 sum_of_angles(
47 encoding_of_inputs(a, b)
48 )
49 )
Listing 8: Mechanistic interpretation of the modular arithmetic model
As noted in Section 5, the discretization operators are necessary for compatibility with Axioms 1 and 2. Without rounding on the first component, we obtain epsilons of 1 for Axioms 1 and 2; epsilons for Axioms 3 and 4 are unaffected by rounding.
For the second component, we apply the equivalence class formulation described in Section 5; however, we strengthen the equivalence criterion by ensuring that equal samples affect neither concrete nor abstract model behavior. This was chosen to avoid the need to sample from the equivalence class on concretization or to select a canonical representative for each class; doing so would have been needed to correctly evaluate Axioms 3 and 4 without such a guarantee. When sampling in this way is feasible, an equivalence relation up to the abstract only model is sufficient. Pseudocode is given in Listing LABEL:lst:equivalence.
1def equivalent(a: cos_sin, b: cos_sin) -> bool:
2 abstract_eq = abstract_components[3](a) == abstract_components[3](b)
3 concrete_eq = concrete_components[3](gammas[2](a)) == concrete_components[3](gammas[2](b))
4 return abstract_eq and concrete_eq
Listing 9: Equivalence relation for second component of mechanistic interpretation of the modular arithmetic model
As for the first component, we obtain epsilons of 1 for Axioms 1 and 2 without applying the equivalence class mapping; by construction, the epsilons for Axioms 3 and 4 are unaffected by the equivalence class operation. Note that simple discretization operators do not suffice: we once again obtain an $\epsilon$ of 1 for Axioms 1 and 2 when rounding to a single decimal place. While strong results with the discretization described above show that the abstract and concrete representations of the output of the second component are equivalent up to their downstream impact on both the concrete and abstract models, results with simpler discretizations indicate that the the concrete representation does not represent the abstract state exactly.
Appendix I A Sketch of Application to Circuits: Indirect Object Identification
As described in Appendix B.1, our axioms may be extended to evaluate circuits-type interpretations, or may be applied directly by linearization of the computational graph of the circuit.
To illustrate the process, we’ll describe how the circuit for Indirect Object Identification (IOI) (Wang et al., 2023) may be evaluated. The IOI circuit consists of five broad categories of heads:
  1. 1.
    Previous token heads, which copy information from the prior token
  2. 2.
    Duplicate token heads, which identify whether there exists any prior duplicate of the current token
  3. 3.
    Induction heads, which serve the same function as duplicate token heads, mediated via the previous token heads
  4. 4.
    S-inhibition heads, which output a signal suppressing attention by name mover heads to duplicated names
  5. 5.
    Name mover heads, which copy names except those suppressed by the S-inhibition heads for output
Figure 18: The IOI circuit, reproduced from Wang et al. (2023).
Note that each of these attention heads computes a well-defined interpretable function which can be represented in our framework. Figure 18 shows the structure of the circuit. We’ll now illustrate the process by which the circuit may be analyzed via the linearization technique from Appendix B.1.2 or the graph axioms from Appendix B.1.1. The circuit in Figure 18 cannot be directly evaluated, as it must be made isomorphic to a decomposition of the concrete model.
To do so, we will first expand the circuit in Figure 18 to a head-by-head graph. Next, for each block with an interpreted head, we construct dummy nodes in the interpretation which correspond to uninterpreted heads; as the interpretation claims that these are uninvolved in the IOI task, these nodes compute the identity function, a constant function or some other no-op which is independent of the IOI task.
Now, we must account for computations performed in uninterpreted blocks and in MLPs. We can express the computation performed by a block of a transformer as a simple computational graph: each interpreted head, and well as the dummy node corresponding to the uninterpreted heads, takes the incoming residual stream as an input and returns an update to the residual stream. These updates are added to the residual stream and processed by the remaining components, such as the MLP and any normalization, which produce the final residual output by the block. For simplicity, we combine any uninterpreted layers into this operation. These nodes, abstractly, update the program state from the potentially redundant input information. In this way, we can construct isomorphic concrete and abstract models from any circuit of this form.
Figure 19: Concrete and abstract models derived from the IOI circuit.
We derive the concrete and abstract models shown in Figure 19. Note that structuring the circuit in this way allows us to clearly define and evaluate a hypothesis on how redundant copies of a given abstract concept are combined downstream. For example, the model might retain independent duplicate copies of the information, one copy per head, retain only the most recently-computed value, or aggregate redundant copies with some weighting.
At this point, we can apply either the linearization technique of Appendix B.1.2 or the generalized axioms of Appendix B.1.1 to evaluate the hypotheses. We do not fully evaluate IOI with our axioms here as Wang et al. (2023) leaves key components of the interpretation unspecified. In particular, the authors do not clearly state how the model combines the information produced by redundant heads performing the same function, and the authors do not conclusively state what the duplicate-token suppressing information output by the S-inhibition heads represents. For example, that information may be represented as the suppressed token itself, via either relative or absolute positions, or a combination of token and position information. A complete analysis with our axioms would enable the analyst to derive clear evidence for the correct hypotheses.